Fermi National Laboratory

Accelerator Update for March 19, 2008 - March 21, 2008


Wednesday March 19

The day shift began with Operations monitoring the Tevatron (TeV) store 5990 (43.44E30) and the Recycler stash (285.3E10), with the Antiproton Source stacking (27.6E10), with the Main Injector (MI), Linac, and Booster operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with NuMI and the SciBooNE / MiniBooNE (BNB) experiments taking beam.

At 8:50 AM, a Linac Klystron RF station (KRF7) tripped off and would not reset. An operator contacted an expert.

Operations put Pbar into standby at 8:54 AM.

At 9:08 AM, Operations received a MI vacuum alarm. Three vacuum pumps (604E, 605A, and 605B) had tripped. They all reset.

At 9:25 AM, a power supply expert reported that the MI-30 quadrupole power supply was getting worse. An expert began his repairs about 45 minutes later.

Operations terminated store 5990 at 9:41 AM.

An expert reported at 9:49 AM, that he had replaced an SRM for KRF7 and it was back online.

Operations reported at 9:53 AM that LCW experts were beginning the TeV LCW fill.

Linac resumed sending beam at 10:20 AM.

Operations reported at 12:15 PM that power was shut off to a leg of Feeder 36 to locate a duct bank. It was necessary to find the duct before a new sewer line gets installed. The power was off to site 50 & 52 from 9:30 to 11:30 AM.

Operations established store 5991 at 1:24 PM with an initial luminosity of 248.58E30.

Operations halted the Linac beam at 1:46 PM so that experts could investigate LRF3 problems.

At 2:44 PM, a Pbar expert reported that a Pbar Debuncher horizontal band TWT had a tunnel problem and its alarm should be bypassed.

Controls experts reported at 3:39 PM that they were rebooting the MI low level RF system. This held off MI beam for about 10 minutes.

Operations reported at 3:59 PM that the Recycler had stashed antiprotons three times during the day shift.

At 5:28 PM, a Booster vacuum pump (IPBD2) tripped off. It reset.

At 6:51 PM, KRF2 tripped off. It reset.

At 9:53 PM, the MI vacuum readbacks stopped reporting. Operations rebooted MIVAC and the readback returned to normal.

At 10 PM, the Linac Klystron vernier tripped off on a LCW flow fault. It reset.

Operations reported at 11:58 PM that the Recycler had stashed five times during the evening shift.

Thursday March 20

The midnight shift began with Operations monitoring store 5991 (84.6E30) and the stash (206.7E10), with Pbar stacking (34.0E10), and with NuMI and BNB taking beam.

Operations terminated store 5991 at 1 AM.

Operations prepared D0 for an access at 1:20 AM. They came out of access at 2:47 AM.

At 3:13 AM, Operations reported that the Pbar autotune program (the Overthruster) used some bad data that caused the P2 transfer line and NuMI beam permit to trip. Operators loaded the program with a good file and everything returned to normal.

Operations put the TeV through a dry squeeze at 3:26 AM.

Operations put the TeV into shot setup at 3:51 AM.

At 4:06 AM, a Booster bulk power supply (B:CEP400) tripped off. It reset.

Operations established store 5993 at 5:40 AM with an initial luminosity of 234.08E30.

Operators had the Pelletron ready for an access by 6:05 AM.

Pbar experts began an AP-10 study at 6:38 AM.

Operators had MI and NuMI prepared for access by 6:40 AM.

Operations reported at 7:54 AM that the Recycler had stashed three times during the midnight shift.

Linac experts reported that they were beginning work on LRF5 at 8:16 AM.

Operators made Pbar ready for access at 8:38 AM.

A water system expert reported at 9:10 AM that FESS personnel would be cleaning the strainers for the MI pond water pumps. The job will take about two and a half hours.

Power supply experts reported at 9:18 AM that the MI D:QF power supply leak had been repaired.

At 9:52 AM, CUB experts reported that they had switched the Booster 95 pump LP2 & LP3 to LP1 & LP2.

Roads and Grounds reported at 10:22 AM that they were conducting a prairie burn in the MI area. The expected it to take about two hours.

Preaccelerator experts made a short access at 11 AM into the I- Source dome.

MI came out of access at 12:18 PM.

Pbar experts came out of their AP0 access at 12:30 PM.

Linac experts reported at 12:41 PM that they had completed their LRF5 studies.

Operations established beam to NuMI at 2:08 PM.

At 2:21 PM, Operations reported some trouble with MIRF11.

Operations established beam to BNB and Pbar at 2:22 PM.

At 2:41 PM, KRF1 tripped off. It reset.

An RF expert investigated MIRF11 and at 2:45 PM brought it back online.

At 2:46 PM, a Booster vacuum pump tripped off and caused BRF13 & 14 to also trip. Everything reset.

Operators reported at 3:08 PM that the Meson MTest area had been searched and secured.

At 3:10 PM, MIRF16 tripped off. It reset.

At 3:43 PM, MI ramp was halted to address MI-10 LCW heat exchanger temperature issues.

The MI ramp was back up at 4:27 PM, cooling was increased to the heat exchanger.

MTest turned on at 4:41 PM.

MTest turned off at 6:34 PM.

Problems with the Linac RF #1 (LRF1) driver caused intermittent beam starting at 10:13 PM and lasted until 11:30 PM. Experts were called in to tune up the driver.

Friday March 21

The midnight shift began with Operations monitoring store 5993(50.58E30) and the stash (90.5E10), with Pbar stacking (29.0E10), and with NuMI and BNB taking beam.

At 12:33 AM, frequent trips of Booster RF #9 (BRF9) resulted in BRF19 being turned on and BRF9 left off.

Store 5993 was terminated at 5:58 AM.

CDF and D0 were prepared for access at 6:13 AM.

Tev A-E and Transfer Hall were prepared for access at 6:52 AM.

At 7:32 AM, Pbar stacking was put into standby for interlock testing in Pre-Vault and Pre-Target.

The Plans for Friday and the Weekend

The plans is to continue running beam to NuMI and Boones. Put a store in the Tev when the collision Hall accesses are complete. To stack, stash and store over the weekend.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/24/2008    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory